OpenStudy (anonymous):

Derivative of 3+xsin(pi(x))

OpenStudy (jdoe0001):

product rule, then chain-rule it

OpenStudy (raden):

you need the product rule if y = uv then y' = u'v +uv'

OpenStudy (anonymous):

I got y′ = (x pi'(x)cos(pi(x)+ sin(pi(x)) but I am not sure if I did it right.

OpenStudy (jdoe0001):

hmmm derivative of x is 1

OpenStudy (jdoe0001):

other than that, looks ok

OpenStudy (anonymous):

wait so what should I change? should I make x=1 instead?

OpenStudy (jdoe0001):

(x pi'(x)cos(pi(x)+ sin(pi(x)) ^ that x

OpenStudy (jdoe0001):

I was wondering about that lone x

OpenStudy (jdoe0001):

\(\bf y=3+xsin(\pi x) \quad y'=[1\cdot sin(\pi x)] +[x\cdot cos(\pi x)\cdot \pi]\)